Suppose that Qd = 200 - 3P and Qs = 20 + 2P. If a tax is imposed on this market such that Qtax = 89, then Pbuy = , Psell = , the size of the tax is , and the deadweight loss is
Answer #1

We are given demand and supply curves and post tax quantity and asked to find price paid by sellers, price paid by buyers, size of tax and deadweight loss

When no tax is imposed, equilibrium is at point where demand is equal to supply.

Equilibrium price is 36 and equilibrium quantity is 92

Producer surplus is 2016 and consumer surplus is 1410.82

After tax is imposed, quantity sold in the market is 89

Price received by sellers is 34.5

Price paid by buyers is 37

Consumer surplus is 1320.315 and producer surplus is 1880.25

Tax revenue is 222.5 and deadweight loss is 3.75

OEMAND CURVE SUPPLY CURVE Qp = 200-3P Qs = 20 +2 p. TAX THAN O O is imOSED Post TAX UNITS ON MARKET SUCH THAT QUANTITY IN MARWILL PRICE, 66.00/2 0377 PB -- --- 1, 6) = - - - - - XE 134.5) Psl--- ---73! > QUANTITY oth 89 92 ss Supply CURVE DDE DEMANAFTER TAX IS IMPOSED OF 2.5 PER UND QUANTITY SOLD IN MARKET is 89 UNITS PRICE PAID BY BUYERS = 37 PRICE PAID BY SELEERS = 34.